# Is the Reversal Curse a Binding Problem? Uncovering Limitations of Transformers from a Basic Generalization Failure

### v1

@@ Abstract
Despite their impressive capabilities, LLMs exhibit a basic generalization failure known as the Reversal Curse, where they struggle to learn reversible factual associations. Understanding why this occurs could help identify weaknesses in current models and advance their generalization and robustness. In this paper, we conjecture that the Reversal Curse in LLMs is a manifestation of the long-standing binding problem in cognitive science, neuroscience and AI. Specifically, we identify two primary causes of the Reversal Curse stemming from transformers’ limitations in conceptual binding: the inconsistency and entanglements of concept representations. We perform a series of experiments that support these conjectures. Our exploration leads to a model design based on JEPA (Joint-Embedding Predictive Architecture) that for the first time breaks the Reversal Curse without side-stepping it with specialized data augmentation or non-causal masking, and moreover, generalization could be further improved by incorporating special memory layers that support disentangled concept representations. We demonstrate that the skill of reversal unlocks a new kind of memory integration that enables models to solve large-scale arithmetic reasoning problems via parametric forward-chaining, outperforming frontier LLMs based on non-parametric memory and prolonged explicit reasoning. [^1] Code and data: https://github.com/OSU-NLP-Group/reversal-curse-binding.
@@ 1 Introduction
There are several pieces of work trying to understand or mitigate the Reversal Curse. Zhu et al. (2024) theoretically shows that reversal cannot be learned for transformers under special settings and assumptions. Lin et al. (2024) shows that the issue may be related to inherent biases in LLMs’ factual recall. Golovneva et al. (2024); Guo et al. (2024a); Lu et al. (2024); Lv et al. (2024); Kitouni et al. (2024) propose specialized data augmentation strategies (e.g., reversing/permuting sentence segments) or non-causal training objectives, which circumvent the problem and reduce generalization demands on models. Overall, existing solutions are rather ad-hoc and fall short of uncovering the potentially more foundational issues behind such a curse—in fact, the very first question still remains a mystery:
Are conventional (autoregressive) transformers fundamentally doomed for learning reversal? [^3] For brevity, “reversal” means “reversal over parametric knowledge” by default. It is well known that LLMs can reverse information when the forward direction is explicitly given in the context. However, this does not lead to any generalizable solution, since it is nearly impossible to prepare in the context all pieces of information that might be needed in reverse, especially when the underlying knowledge required for the query is massive and layered.
@@
Surprisingly, the answer is “No”. Our first major finding is that standard transformers can learn reversal without any specialized data augmentation or modifications to the architecture or objective, when the inputs are represented and perceived at the abstract concept level (Figure 1(a)). We then focus on the gap between abstract and real settings, where inputs are instead at the surface form level. Our investigations lead us to hypothesize that the Reversal Curse is fundamentally a manifestation of the long-standing binding problem in cognitive science, neuroscience and AI, which is concerned with the mechanisms for natural or artificial neural networks to combine information distributed throughout the network to form integrated percepts and knowledge (Roskies, 1999; Engel & Singer, 2001; Zimmer et al., 2006; Greff et al., 2020). Specifically, we conjecture that the Reversal Curse is primarily caused by two limitations of conceptual binding in transformers, the inconsistency and entanglements of concept representations:
  • Inconsistency. While many existing studies show that LLMs could form internal concepts and even “world models” from surface-level predictions (Meng et al., 2022; Geva et al., 2023; Lad et al., 2024; Kaplan et al., 2025; Li et al., 2023; Gurnee & Tegmark, 2024), we believe they are still unable to adequately learn consistent concept representations across various places within the network under different contexts. Specific to reversal, we conjecture that transformers fail to bind representations of the same underlying entity when it switches roles between perceived subjects and predicted objects (Figure 1(b), Left), which makes the model’s acquired knowledge fragmented and impedes the learning of reversal.
@@
A series of quantitative experiments support our hypotheses, and inform two designs for mitigating the Reversal Curse: 1) performing autoregressive prediction at the concept level, akin to Joint-Embedding Predictive Architectures (JEPA) (LeCun, 2022) and concept models (Barrault et al., 2024), and 2) building dedicated recognition modules which support disentangled concept representations. We show that 1) a model design based on JEPA and in-batch contrastive learning could, for the first time to our knowledge, break the Reversal Curse with non-trivial performance without circumventing the problem, but suffers from entanglements that scale with model depth; 2) incorporating special memory layers (Sukhbaatar et al., 2015; Berges et al., 2024) into the recognition module could further boost generalization.
Our high-level goal is to improve the parametric memory of AI models, which we believe is important for handling difficult knowledge and reasoning tasks. Towards this end, we demonstrate that the reversal skill unlocks a new kind of parametric memory integration, that allows models to implicitly “chain” pieces of information multiple steps away. This enables the model to perform parametric forward-chaining during information internalization to solve large-scale arithmetic reasoning problems with impressive performance, outperforming frontier LLMs based on non-parametric memory.
To summarize, our work 1) contributes towards more fundamentally understanding and addressing the Reversal Curse, and more importantly, 2) connects the Reversal Curse with the more foundational problem of improving conceptual binding and generalization in AI models, rigorously establishing the concrete challenges for the broader research community.
@@ 2 Learning Reversal at the Concept Level
Humans think and learn at the concept level. When reading a sentence, we (usually subconsciously) parse and map the words into concepts, and update the concept representations and associations upon encountering new information (Collins & Loftus, 1975; Jackendoff, 1995). While transformers fail to learn reversal in real settings, do they first have appropriate inductive biases to learn reversal at the abstract concept level?
@@
Setup. We prepare a set of relation pairs $\{(r_{i},r_{i}^{-1})\mid i=1,\dots,N\}$, and two disjoint sets of entities $\mathcal{E}_{A}$ (for learning) and $\mathcal{E}_{B}$ (for testing). We focus on one-to-one relations that give a unique object entity for each fact. We synthesize facts separately over $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ by randomly pairing the entities for each $(r_{i},r_{i}^{-1})$, and form a pair of facts which are reverses of each other over each entity pair. Through this, we obtain two sets of facts $D_{A}$ and $D_{B}$ over $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ respectively. The training data contains all of $D_{A}$ (for the model to induce the rules) and one random direction from each pair of facts in $D_{B}$, where its reverse goes into the test set. We set $N=6$, and vary $|\mathcal{E}_{A}|$ while keeping the ratio $|\mathcal{E}_{A}|:|\mathcal{E}_{B}|=5:3$. Importantly, each concept (entity/relation) is directly represented by its own learnable embedding, without attaching to surface-level names. We train standard decoder-only transformers as in GPT-2 (Radford et al., 2019) (with $768$ hidden dimensions and $12$ attention heads) to predict the object entity in each fact, with cross-entropy loss over embeddings of all concepts. [^4] Positional encodings do not affect the results from preliminary experiments. We train models for a large number of steps ($63e$ 6) and report the highest mean reciprocal rank (MRR) on the test examples achieved among different model checkpoints. More training details are included in Appendix A.
@@
#Layer $=1$ $0.823$ $0.861$ $0.947$ $0.964$
#Layer $=6$ $0.890$ $0.858$ $0.951$ $0.861$
#Layer $=12$ $0.810$ $0.878$ $0.951$ $0.960$
#Layer $=18$ $0.823$ $0.850$ $0.944$ $0.975$
Table 1: Mean reciprocal rank (MRR) achieved by standard transformers in the abstract setting, where inputs are represented at the concept level.
Transformers can learn reversal at the concept level. Results are in Table 1. Surprisingly, in contrast with the negative results and views in previous studies, we find that transformers can learn reversal with high performance without any specialized training objectives or data augmentation. Models with different depths could all strongly generalize, where the performance overall increases with $|\mathcal{E}_{A}|$. [^5] Note that with a larger $|\mathcal{E}_{A}|$, while there are more training data which benefits learning, it is also harder to achieve higher performance since the prediction is contrasted with a larger population. Despite being extremely straightforward, this positive result becomes one of the major findings in this work. The critical question then arises: If transformers can learn reversal at the abstract concept level, why do they fail in realistic settings?
@@ 3 The Binding Problem Underlying Surface-level Predictions
The binding problem. Our central thesis is that the Reversal Curse is a manifestation of the long-standing binding problem in cognitive science, neuroscience and AI, which is concerned with the mechanisms for natural or artificial neural networks to bind information distributed throughout the network and form integrated percepts and knowledge (Roskies, 1999; Engel & Singer, 2001; Zimmer et al., 2006; Greff et al., 2020). The binding problem could be divided into two major types: perceptual binding and conceptual binding. Perceptual binding refers to the combination of features/attributes from raw inputs into cohesive concepts, typically occurring during low-level perception/recognition (Von Der Malsburg, 1994; Tallon-Baudry & Bertrand, 1999; Singer, 2007; Palmigiano et al., 2017). Conceptual binding is centered around forming unified and integrated long-term conceptual knowledge, which occurs during high-level semantic processing and memory consolidation (McNorgan et al., 2011; Opitz, 2010; Murre et al., 2006; Patterson et al., 2007; Ralph et al., 2017).
Numerous studies show that LLMs have no trouble learning perceptual binding through surface-level predictions. For example, prior research finds that LLMs usually perform “detokenization” and combine sub-word tokens into cohesive concept representations at the end of surface names within early layers (Meng et al., 2022; Geva et al., 2023; Lad et al., 2024; Yang et al., 2024; Kaplan et al., 2025); in upper layers, tokens of the output surface name beyond the immediate next token are usually (often-times, linearly) encoded in the hidden states (Pal et al., 2023; Belrose et al., 2023; Wu et al., 2024; Cai et al., 2024). In addition, studies tracking the evolvement of LLMs’ internal states during inference suggest that they also “think” in an abstract concept space on top of perceptual binding within the middle layers (Geva et al., 2022; Wendler et al., 2024; Lad et al., 2024; Sun et al., 2025). An illustration is in Figure 1(b). These findings indicate that the issues lie not in perceptual binding, but in the specific representations learned under surface-level prediction. Upon closer examination, we identify two key potential factors contributing to the failure in learning reversal, both stemming from transformers’ deficiency in conceptual binding: inconsistency and entanglements.
@@ 3.1 Inconsistency of Concept Representations
Conceptual consistency is challenging to learn well within the design of current transformers even with an abundance of data, due to the dynamic and open-ended nature of concepts. Firstly, concepts in transformers could emerge at various locations/subspaces, necessitating a dynamic tracking and routing mechanism. Secondly, concepts are continuously assimilated instead of coming from a fixed vocabulary, which requires a systematic design that can establish the binding of newly acquired concepts automatically. This level of systematicity is not achieved even for the much simpler problem of binding tokens at the input/output levels, especially for models with untied input/output embeddings (which is common in most LLMs today)–—when new tokens are introduced to the vocabulary, transformers still have to rely on dedicated data to establish their binding.
@@ 3.2 Entanglements of Concept Representations
To illustrate, consider the last MLP layer in the recognition module before concept representations are formed, as shown in Figure 1(b), Right. Suppose we have two activated concepts $a$ and $b$ in the current learning step, which have MLP hidden activations $\alpha,\beta$ respectively. Let the output projection matrix be $V$ where $v_{i}$ is its $i$-th column, and hence $a=\sum_{i}\alpha_{i}v_{i},b=\sum_{i}\beta_{i}v_{i}$. [^7] Here we ignore the residual connection and bias terms for simplicity. During learning when the loss is $L$, the negative gradients $-\partial L/\partial a$ and $-\partial L/\partial b$ represent the “desired directions” for updating $a$ and $b$. Assuming for now that $\alpha,\beta$ remain constant, we could compute the updates of concept representations $\Delta a$ and $\Delta b$ after a gradient descent step with step size $\eta$:
@@
$\Delta a=\sum_{i}\alpha_{i}(v_{i}-\eta\frac{\partial L}{\partial v_{i}})-a=-\eta||\alpha||^{2}\frac{\partial L}{\partial a}-\eta\alpha^{T}\beta\frac{\partial L}{\partial b},$
$\Delta b=\sum_{i}\beta_{i}(v_{i}-\eta\frac{\partial L}{\partial v_{i}})-b=-\eta||\beta||^{2}\frac{\partial L}{\partial b}-\eta\alpha^{T}\beta\frac{\partial L}{\partial a}.$
We can see that each concept is not updated in the direction of its negative gradient; rather, the updates are mixed with gradients from other concepts, where the level of entanglements is decided by $\alpha^{T}\beta$, i.e., how strong the hidden activations of $a$ and $b$ overlap (red neurons in Figure 1(b), Right). [^8] These entanglements clearly extend to momentum-based updates in modern optimizers. This overlap is in turn determined by the surface form names of $a,b$ and the configuration of lower-level weights, which are also subject to change and could add further complications. This is very troublesome, especially given that concept names could be almost arbitrary and exhibit all kinds of correlations. For example, imagine two different people with somewhat similar names. While there are pattern overlaps during recognition, after it is complete, they should become two distinct objects, and the information that we wish to store on each should be stored independently and not interfere. However, as we can see, for transformers with gradient-based optimization, the overlaps in activation patterns effectively cause the learning of different concepts to “mix”, which could adversely influence the training dynamics and generalization.
@@ 3.3 Experiments
Transformers fail to learn reversal with surface-level predictions. We first conduct a series of experiments on models trained with standard language modeling objectives at the surface level. All model variants completely fail to generalize. Specifically, we first find that vanilla transformers with different depths do not generalize. Since this failure could stem from a lack of perceptual binding within the model, we implement the known structures in LLMs as illustrated in Figure 1(b) by explicitly building into the model a recognition module that maps each concept name to a single hidden state, and a verbalization module which decodes the predicted hidden state back into concept names. Again, we find no signs of generalization. We also test both tied/untied embeddings across all variants above, and observe no difference. Overall, transformers with surface-level predictions completely fail to learn reversal. These results mostly serve as validations and should come with little surprise, since LLMs trained on much more diverse and extensive corpora exhibit the Reversal Curse.
@@
We begin with explicitly encouraging conceptual consistency in reversal. One key observation is that the well-known Joint-Embedding Predictive Architecture (JEPA) (LeCun, 2022), which conducts predictions at the abstract representation space instead of raw input space, could perfectly serve this purpose if the abstract representations for prediction are at the concept level. This is somewhat a “coincidence” since the original motivation of JEPA is to ignore unpredictable/unimportant information in the inputs, which is not related to the focus here. We experiment with a simple instantiation of JEPA based on in-batch contrastive learning. Specifically, autoregressive prediction is now done at the concept level encoded by the recognition module, where the representations of other “activated” concepts besides the ground truth within the same batch serve as negatives for standard contrastive loss (more details in Appendix B). We evaluate the quality of learned representations by comparing the predicted state against representations of all concepts. We experiment with different configurations of the model in terms of the depth of the recognition module (#Rec) and the depth of the semantic module (#Sem), the semantic processing component on top of the recognition module.
JEPA unlocks generalization, but suffers from entanglements. Results are in Figure 2 (Left). By simply encouraging conceptual consistency (with JEPA), the model can achieve highly non-trivial generalization. To our knowledge, this is the first-ever model design that breaks the reversal curse without side-stepping the core problem. Another important observation is that the performance decreases when the model becomes deeper. Specifically, generalization consistently worsens when increasing either 1) the depth ratio between the semantic and recognition module, or 2) the overall depth of the model while keeping the ratio fixed. This strongly indicates that the effect of entanglements scales with model depth. This is intuitive since the representational distortions caused by entanglements accumulate throughout the layers. We also examine the impact of multiplicity, where the results are in Figure 2 (Right). It could be seen that increasing the multiplicity severely hurts generalization, especially with deeper models whose performance could drop to near zero with a mere multiplicity of $20$. Overall, the results here suggest that current models likely learn a low degree of conceptual consistency, and even if not, it is still challenging for them to learn reversal due to the effects of entanglements.
@@
Results are in Figure 3. It can be observed that increasing the width does aid generalization, but only incrementally (Figure 3, Left). On the other hand, the specially designed memory layer could significantly enhance performance, though generalization still mildly declines with more semantic layers (Figure 3, Right). These results also confirm that the limited generalization observed with standard transformer layers as the recognition module is not due to insufficient capacity, since the module with two $1280$-width transformer layers already has nearly the same amount of effective parameters as the memory layer ($61.1$M vs. $61.4$M). Overall, these results underscore the importance of thoughtful designs in addressing the issue of entanglements apart from scaling. Our findings also provide a concrete example that memory layers can improve generalization, corroborating recent efforts on scaling memory layers that report enhanced performance on general-domain tasks (Berges et al., 2024).
4 Parametric Forward-Chaining for Large-Scale Arithmetic Reasoning

### v2

@@ Abstract
Despite their impressive capabilities, LLMs exhibit a basic generalization failure known as the Reversal Curse, where they struggle to learn reversible factual associations. Understanding why this occurs could help identify weaknesses in current models and advance their generalization and robustness. In this paper, we conjecture that the Reversal Curse in LLMs is a manifestation of the long-standing binding problem in cognitive science, neuroscience and AI. Specifically, we hypothesize two primary causes of the Reversal Curse stemming from transformers’ limitations in conceptual binding: the inconsistency and entanglements of concept representations. We perform a series of experiments that support these conjectures. Our exploration leads to a model design based on JEPA (Joint-Embedding Predictive Architecture) that for the first time breaks the Reversal Curse without side-stepping it with specialized data augmentation or non-causal masking, and moreover, generalization could be further improved by incorporating special memory layers that support disentangled concept representations. Our research opens up the broader fundamental challenge of designing models capable of learning systematic conceptual binding with less human scaffolding. [^1] Code and data: https://github.com/OSU-NLP-Group/reversal-curse-binding.
@@ 1 Introduction
There are several pieces of work trying to understand or mitigate the Reversal Curse. Zhu et al. (2024) theoretically shows that reversal cannot be learned for transformers under special settings and assumptions. Lin et al. (2024) shows that the issue may be related to inherent biases in LLMs’ factual recall. Golovneva et al. (2024); Guo et al. (2024); Lu et al. (2024); Lv et al. (2024); Kitouni et al. (2024) propose specialized data augmentation strategies (e.g., reversing/permuting sentence segments) or non-causal training objectives, which circumvent the problem and reduce generalization demands on models. Overall, existing solutions are ad-hoc and fall short of uncovering the potentially more foundational issues behind such a curse—in fact, the very first question still remains a mystery:
Are conventional (autoregressive) transformers fundamentally doomed for learning reversal?
@@
Surprisingly, the answer is “No”. Our first major finding is that standard transformers can learn reversal without any specialized data augmentation or modifications to the architecture or objective, when the inputs are represented and perceived at the abstract concept level (Figure 1(a)). We then focus on the gap between abstract and real settings, where inputs are instead at the surface form level. Our investigations lead us to hypothesize that the Reversal Curse is fundamentally a manifestation of the long-standing binding problem in cognitive science, neuroscience and AI, which is concerned with the mechanisms for natural or artificial neural networks to combine information distributed throughout the network to form integrated percepts and knowledge (Roskies, 1999; Engel and Singer, 2001; Zimmer et al., 2006; Greff et al., 2020). Specifically, we conjecture that the Reversal Curse is primarily caused by two limitations of conceptual binding in transformers, the inconsistency and entanglements of concept representations:
  • Inconsistency. While many existing studies show that LLMs could form internal concepts and even “world models” from surface-level predictions (Meng et al., 2022; Geva et al., 2023; Lad et al., 2024; Kaplan et al., 2025; Li et al., 2023; Gurnee and Tegmark, 2024), we hypothesize that they are still unable to adequately learn consistent concept representations across various places within the network under different contexts. Specific to reversal, we conjecture that transformers fail to bind representations of the same underlying entity when it switches roles between perceived subjects and predicted objects (Figure 1(b), Left), which makes the model’s acquired knowledge fragmented and impedes the learning of reversal.
@@
A series of quantitative experiments support our hypotheses, and inform two novel designs for mitigating the Reversal Curse: 1) performing autoregressive prediction at the concept level, akin to Joint-Embedding Predictive Architectures (JEPA) (LeCun, 2022) and concept models (Barrault et al., 2024), and 2) building dedicated recognition modules which support disentangled concept representations. We show that 1) a model design based on JEPA and in-batch contrastive learning could, for the first time to our knowledge, break the Reversal Curse with non-trivial performance without circumventing the problem, but suffers from entanglements that scale with model depth; 2) incorporating special memory layers (Sukhbaatar et al., 2015; Berges et al., 2024) into the recognition module could further boost generalization. We also demonstrate that the reversal skill unlocks a new kind of parametric memory integration, which allows models to perform parametric forward-chaining during information internalization to solve large-scale arithmetic reasoning problems with impressive performance, outperforming frontier LLMs based on non-parametric memory.
To summarize, our work 1) contributes towards understanding and addressing the Reversal Curse, and more importantly, 2) connects the Reversal Curse with the more foundational problem of improving conceptual binding and generalization in AI models, rigorously establishing the concrete challenges for the broader research community.
@@ 2 Learning Reversal at the Concept Level
Humans think and learn at the concept level. When reading a sentence, we (usually subconsciously) parse and map the words into concepts, and update the concept representations and associations upon encountering new information (Collins and Loftus, 1975; Jackendoff, 1995). While transformers fail to learn reversal in real settings, do they first have appropriate inductive biases to learn reversal at the abstract concept level?
@@
Setup. We prepare a set of relation pairs $\{(r_{i},r_{i}^{-1})\mid i=1,\dots,N\}$, and two disjoint sets of entities $\mathcal{E}_{A}$ (for learning) and $\mathcal{E}_{B}$ (for testing). We focus on one-to-one relations that give a unique object entity for each fact. We synthesize facts separately over $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ by randomly pairing the entities for each $(r_{i},r_{i}^{-1})$, and form a pair of facts which are reverses of each other over each entity pair. Through this, we obtain two sets of facts $D_{A}$ and $D_{B}$ over $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ respectively. The training data contains all of $D_{A}$ (for the model to induce the rules) and one random direction from each pair of facts in $D_{B}$, where its reverse goes into the test set. We set $N=6$, and vary $|\mathcal{E}_{A}|$ while keeping the ratio $|\mathcal{E}_{A}|:|\mathcal{E}_{B}|=5:3$. Importantly, each concept (entity/relation) is directly represented by its own learnable embedding, without attaching to surface-level names. We train standard decoder-only transformers as in GPT-2 (Radford et al., 2019) (with $768$ hidden dimensions and $12$ attention heads) to predict the object entity in each fact, with cross-entropy loss over embeddings of all concepts. We train models for a large number of steps ($3e6$) and report the highest mean reciprocal rank (MRR) on the test examples achieved among different model checkpoints. More training details are included in Appendix A.
@@
#Layer =1=1 $0.823$ $0.861$ $0.947$ $0.964$
#Layer =6=6 $0.890$ $0.858$ $0.951$ $0.861$
#Layer =12=12 $0.810$ $0.878$ $0.951$ $0.960$
#Layer =18=18 $0.823$ $0.850$ $0.944$ $0.975$
Table 1: Mean reciprocal rank (MRR) achieved by standard transformers in the abstract setting, where inputs are represented at the concept level. Here $\mathcal{E}_{A}$ is the set of entities in the training set for learning/inducing the rules of reversal.
Transformers can learn reversal at the concept level. Results are in Table 1. Surprisingly, in contrast with the negative results and views in previous studies, we find that transformers can learn reversal with high performance without any specialized training objectives or data augmentation. Models with different depths could all strongly generalize, where the performance overall increases with $|\mathcal{E}_{A}|$. Despite being extremely straightforward, this positive result becomes one of the major findings in this work. The critical question then arises: If transformers can learn reversal at the abstract concept level, why do they fail in realistic settings?
@@ 3 The Binding Problem Underlying Surface-level Predictions
The binding problem. Our central thesis is that the Reversal Curse is a manifestation of the long-standing binding problem in cognitive science, neuroscience and AI, which is concerned with the mechanisms for natural or artificial neural networks to bind information distributed throughout the network and form integrated percepts and knowledge (Roskies, 1999; Engel and Singer, 2001; Zimmer et al., 2006; Greff et al., 2020). The binding problem could be divided into two major types: perceptual binding and conceptual binding. Perceptual binding refers to the combination of features from raw inputs into cohesive concepts, typically occurring during low-level perception/recognition (Von Der Malsburg, 1994; Tallon-Baudry and Bertrand, 1999; Singer, 2007; Palmigiano et al., 2017). Conceptual binding is centered around forming unified and integrated long-term knowledge, which occurs during high-level semantic processing and memory consolidation (McNorgan et al., 2011; Opitz, 2010; Murre et al., 2006; Patterson et al., 2007; Ralph et al., 2017).
Numerous studies show that LLMs have no trouble learning perceptual binding through surface-level predictions. For instance, prior work finds that detokenization is typically carried out in a “recognition module” within the lower layers, where subword tokens are combined and mapped into cohesive concept representations at the end of surface names (Meng et al., 2022; Geva et al., 2023; Lad et al., 2024; Yang et al., 2024; Kaplan et al., 2025); in upper layers, tokens of the output surface name beyond the immediate next token are also usually (often-times, linearly) encoded in the hidden states (Pal et al., 2023; Belrose et al., 2023; Wu et al., 2024; Cai et al., 2024). Complementary studies tracing the evolvement of LLMs’ internal states during inference suggest that, beyond perceptual binding, they also “think” in an abstract concept space within a “semantic module” in the middle layers (Geva et al., 2022; Wendler et al., 2024; Lad et al., 2024; Sun et al., 2025). An illustration is in Figure 1(b). These findings indicate that the issues lie not in perceptual binding, but in the specific representations learned under surface-level prediction. Upon closer examination, we identify two key potential factors contributing to the failure in learning reversal, both stemming from transformers’ deficiency in conceptual binding: inconsistency and entanglements.
@@ 3.1 Inconsistency of Concept Representations
Conceptual consistency is challenging to learn well within the design of current transformers even with an abundance of data, due to the dynamic and open-ended nature of concepts. Firstly, concepts in transformers could emerge at various locations/subspaces, necessitating a mechanism for dynamically tracking and connecting the concept representations across different regions of the network. Secondly, concepts are continuously assimilated instead of coming from a fixed vocabulary, which requires a systematic design that can establish the binding of newly acquired concepts automatically. This level of systematicity is not achieved even for the much simpler problem of binding tokens at the input/output levels, especially for models with untied input/output embeddings (which is common in most LLMs today)–—when new tokens are introduced to the vocabulary, transformers still have to rely on dedicated data to establish their binding.
@@ 3.2 Entanglements of Concept Representations
To illustrate, consider the last MLP layer in the recognition module before concept representations are formed, as shown in Figure 1(b), Right. Suppose we have two activated concepts $a$ and $b$ in the current learning step, which have MLP hidden activations $\alpha,\beta$ respectively. Let the output projection matrix be $V$ where $v_{i}$ is its $i$-th column, and hence $a=\sum_{i}\alpha_{i}v_{i},b=\sum_{i}\beta_{i}v_{i}$. [^4] Here we ignore the residual connection and bias terms for simplicity. During learning when the loss is $L$, the negative gradients $\partial L/\partial a$ and $\partial L/\partial b$ represent the “desired directions” for updating $a$ and $b$. Assuming for now that $\alpha,\beta$ remain constant, we could compute the updates of concept representations $\Delta a$ and $\Delta b$ after a gradient descent step with step size $\eta$:
@@
[…]
We can see that each concept is not updated in the direction of its negative gradient; rather, the updates are mixed with gradients from other concepts, where the level of entanglements is decided by $\alpha^{T}\beta$, i.e., how strong the hidden activations of $a$ and $b$ overlap (red neurons in Figure 1(b), Right). [^5] These entanglements clearly extend to momentum-based updates in modern optimizers. This overlap is in turn determined by the surface form names of $a,b$ and the configuration of lower-level weights, which are also subject to change and could add further complications. Note that while our analysis here focuses on one specific layer, the effect of entanglements naturally scales with model depth as the representational distortions accumulate throughout layers. This is very problematic, especially given that concept names could be almost arbitrary and exhibit all kinds of correlations. For example, imagine two different people with somewhat similar names. While there are pattern overlaps during recognition, after it is complete, they should become two distinct objects, and the information that we wish to store on each should be stored independently and not interfere. However, as we can see, for transformers with gradient-based optimization, the overlaps in activation patterns effectively cause the learning of different concepts to “mix”, which could adversely influence the training dynamics and generalization.
@@ 3.3 Experiments
We first conduct a series of experiments on models trained with standard language modeling objectives at the surface level, and confirm that transformers fail entirely (0.0% accuracy) to learn reversal from surface-level predictions, regardless of architectural variants (e.g., depth, tied vs. untied embeddings). These findings align with prior work (Golovneva et al., 2024; Allen-Zhu, 2024).
@@
Figure 2: Illustration of JEPA with in-batch contrastive learning.
@@
We begin with explicitly encouraging conceptual consistency in reversal. One key observation is that the well-known Joint-Embedding Predictive Architecture (JEPA) (LeCun, 2022), which conducts predictions at the abstract representation space instead of raw input space, could perfectly serve this purpose if the abstract representations for prediction are at the concept level. This is somewhat a “coincidence” since the original motivation of JEPA is to ignore unpredictable/unimportant information in the inputs, which is not related to the focus here. We experiment with a simple instantiation of JEPA based on in-batch contrastive learning, illustrated in Figure 2. Here, autoregressive prediction is done at the concept level encoded by the recognition module, where the representations of other “activated” concepts within the same batch (besides the ground truth) serve as negatives for the standard InfoNCE contrastive loss (van den Oord et al., 2019) (additional details in Appendix B). We evaluate the quality of learned representations by comparing the predicted state against representations of all concepts. We experiment with different configurations of the model in terms of the depth of the recognition module (#Rec) and the depth of the semantic module (#Sem), the semantic processing component on top of the recognition module.
JEPA unlocks generalization, but suffers from entanglements. Results are in Figure 3 (Left). By simply encouraging conceptual consistency (with JEPA), the model can achieve highly non-trivial generalization. To our knowledge, this is the first-ever model design that breaks the reversal curse without side-stepping the core problem. Another important observation is that the performance decreases when the model becomes deeper. Specifically, generalization consistently worsens when increasing either 1) the depth ratio between the semantic and recognition module, or 2) the overall depth of the model while keeping the ratio fixed. This strongly corroborates our analysis that the effect of entanglements scales with model depth. We also examine the impact of multiplicity, where the results are in Figure 3 (Right). It could be seen that increasing the multiplicity severely hurts generalization, especially with deeper models whose performance could drop to near zero with a mere multiplicity of $20$. Overall, the results here suggest that current models likely learn a low degree of conceptual consistency, and even if not, it is still challenging for them to learn reversal due to the effects of entanglements.
@@
It can be observed that increasing the width does aid generalization, but only incrementally (Figure 4, Left). Meanwhile, the specially designed memory layer could significantly enhance performance, though generalization still mildly declines with more semantic layers (Figure 4, Right). These results also confirm that the limited generalization observed with standard transformer layers as the recognition module is not due to insufficient capacity, since the module with two $1280$-width transformer layers already has nearly the same amount of effective parameters as the memory layer (61.161.1M vs. 61.461.4M). We further validate that standard transformers equipped with memory layers also fail to generalize, confirming that the inconsistency inherent under surface-level prediction remains a primary blocker; memory layers are effective only when this inconsistency is resolved. Overall, these results underscore the importance of thoughtful designs in addressing the issue of entanglements apart from scaling. Our findings also provide a concrete example that memory layers can improve generalization, corroborating recent efforts on scaling memory layers that report enhanced performance on general-domain tasks (Berges et al., 2024).
4 Arithmetic Reasoning via Parametric Forward-Chaining